# On higher-dimensional symmetric designs

### v1

@@ Abstract.
We study two kinds of generalizations of symmetric block designs to higher dimensions, the so-called $\mathcal{C}$-cubes and $\mathcal{P}$-cubes. For small parameters all examples up to equivalence are determined by computer calculations. Known properties of automorphisms of symmetric designs are extended to autotopies of $\mathcal{P}$-cubes, while counterexamples are found for $\mathcal{C}$-cubes. An algorithm for the classification of $\mathcal{P}$-cubes with prescribed autotopy groups is developed and used to construct more examples. A linear bound on the dimension of difference sets for $\mathcal{P}$-cubes is proved and shown to be tight in elementary abelian groups. The construction is generalized to arbitrary groups by introducing regular sets of (anti)automorphisms.
@@ 1. Introduction
Here $I$ is the identity matrix and $J$ is the all-ones matrix. An $n$-cube of $(v,k,\lambda)$ designs [21] is an $n$-dimensional $v\times\cdots\times v$ matrix such that its every $2$-section is an incidence matrix of a $(v,k,\lambda)$ design. Sections of dimension $2$ are submatrices obtained by fixing all but two coordinates. This generalization is a special case of W. de Launey’s proper $n$-dimensional transposable designs; see [5, Definitions 2.1, 2.6, 2.7, and Example 2.2]. The set of all $n$-cubes of $(v,k,\lambda)$ designs was denoted by $\mathcal{C}^{n}(v,k,\lambda)$ in [21]. We shall refer to them as $\mathcal{C}$-cubes.
The second generalization was introduced in [22] under the name of $(v,k,\lambda)$ projection $n$-cubes. These are $n$-dimensional $v\times\cdots\times v$ matrices with $\{0,1\}$-entries such that every $2$-projection is an incidence matrix of a $(v,k,\lambda)$ design. If $C$ is an $n$-dimensional matrix and $1\leq x<y\leq n$, the projection $\Pi_{xy}(C)$ is the $2$-dimensional matrix with $(i_{x},i_{y})$-entry
@@
The sum is taken over all $n$-tuples $(i_{1},\ldots,i_{n})\in\{1,\ldots,v\}^{n}$ with fixed coordinates $i_{x}$ and $i_{y}$ in a field of characteristic 00. This definition was inspired by Room squares [10], which are generalized to $n$-dimensional Room cubes in an analogous way. In [22], the set of all $(v,k,\lambda)$ projection $n$-cubes was denoted by $\mathcal{P}^{n}(v,k,\lambda)$. We shall refer to them as $\mathcal{P}$-cubes.
The purpose of this paper is to study both types of cubes and to compare their properties. For dimension $n=2$ both are just incidence matrices of symmetric designs, but for $n\geq 3$ there are significant differences. For example, it was shown in [22, Theorem 2.8] that the dimension of $(v,k,\lambda)$ projection cubes with $k\geq 2$ is bounded by
$n\leq\frac{v(v+1)}{2}.$ (2)
@@
The layout of our paper is as follows. In Section 2 we recall the definitions of isotopy and equivalence of $n$-dimensional incidence cubes. We discuss the numbers of inequivalent cubes in $\mathcal{C}^{n}(v,k,\lambda)$ and $\mathcal{P}^{n}(v,k,\lambda)$ for $k=1$ and $k=2$. The main result in this section is Teorem 2.4, giving a complete enumeration up to equivalence of $\mathcal{P}$-cubes for parameters $(7,3,1)$, $(7,4,2)$ and all dimensions $n$. The proof is a computer calculation based on an algorithm that successively increases the dimension. The largest possible dimensions are $\nu(7,3,1)=7$ and $\nu(7,4,2)=9$.
@@
In Section 4 we study $\mathcal{P}^{n}(v,k,\lambda)$-cubes constructed from higher-dimensional difference sets. Theorem 4.1 gives a much sharper bound than (2) on the dimension: $n\leq v$. Building on results from [22], we compute new values of $\mu_{G}(v,k,\lambda)$, the largest dimension of $(v,k,\lambda)$ difference sets in the group $G$. For elementary abelian groups $G$, Theorem 4.4 shows that the bound is tight, i.e. $\mu_{G}(v,k,\lambda)=v$ holds whenever difference sets exist. Theorem 4.7 generalizes the construction to arbitrary groups $G$ based on the notion of regular sets of (anti)automorphisms. A nice consequence is Corollary 4.8, showing that cyclic $(v,k,\lambda)$ difference sets extend at least to dimension $p$, where $p$ is the smallest prime divisor of $v$.
In the final Section 5 we put forward some observations based on data in Table 4. The table contains numbers of inequivalent $\mathcal{P}^{n}(v,k,\lambda)$-cubes constructed from difference sets. We have only partial explanations for apparent symmetries of the numbers and hope that the observations could lead to new theorems.
Some of our results have traditional formal proofs, while others are proved by computer calculations. There are numerous connections between the two types of results in both directions. In Section 3, formal results about autotopies enable computer classifications of $\mathcal{P}$-cubes with prescribed autotopy groups. In Section 4, computer classifications of higher-dimensional difference sets provide impetus for formal results, notably Theorems 4.4 and 4.7. Higher-dimensional incidence cubes are most easily explored by studying examples and performing experiments on a computer. Tools for examining $\mathcal{C}$- and $\mathcal{P}$-cubes of symmetric designs are available in the GAP package Prescribed automorphism groups [12, 20]. Plenty of examples are given in this paper, as well as in [21, 22].
@@ 2. Equivalence and classification
The classification problem is to determine the equivalence classes of $\mathcal{C}$- and $\mathcal{P}$-cubes for given parameters $(v,k,\lambda)$ and dimension $n$. We start with the degenerate case $k=1$. Incidence matrices of symmetric $(v,1,0)$ designs are permutation matrices of order $v$. They are all equivalent, since the rows and columns can be permuted to get the identity matrix $I$. This extends straightforwardly to $\mathcal{P}^{n}(v,1,0)$-cubes, which are all equivalent with
@@
On the other hand, the classification of $\mathcal{C}^{n}(v,1,0)$-cubes is a difficult problem. For $n=3$ they are in $1$-to-$1$ correspondence with latin squares $L=(\ell_{i_{1}i_{2}})$ of order $v$ by $C(i_{1},i_{2},i_{3})=[\ell_{i_{1}i_{2}}=i_{3}]$. The square bracket $[P]$ is the Iverson symbol, taking the value $1$ if $P$ is true, and 00 otherwise [19]. The number of main classes of latin squares has been determined by computer calculations up to $v=11$ [13, 25]. Cubes in $\mathcal{C}^{n}(v,1,0)$ of arbitrary dimension $n$ are in $1$-to-$1$ correspondence with latin hypercubes of order $v$ and dimension $n-1$. There have been several definitions of latin hypercubes in the literature; the suitable definition for our purpose is used in [27]. There, classification is performed for $n=4$, $v\leq 6$ and $n=5,6$, $v\leq 5$.
@@
Cubes in $\mathcal{P}^{n}(3,2,1)$ have been classified directly in [22]. The result is reproduced in the first row of Table 1. We see that $\nu(3,2,1)=5$; the bound (2) gives $\nu(3,2,1)\leq 6$ and is not tight. The next feasible parameters are $(7,3,1)$. It is well known that the Fano plane is the unique $(7,3,1)$ design. In [22], two inequivalent $\mathcal{P}^{3}(7,3,1)$-cubes are presented as Examples 2.2 and 2.6. Here we present a full classification of $\mathcal{P}^{n}(7,3,1)$ and $\mathcal{P}^{n}(7,4,2)$-cubes.
@@
The proof is a computer calculation that was carried out using the orthogonal array representation of $\mathcal{P}$-cubes. A function $C:\{1,\ldots,v\}^{n}\to\{0,1\}$ is the characteristic function of a set of $n$-tuples
@@
If $C\in\mathcal{P}^{n}(v,k,\lambda)$, then $\overline{C}$ is an orthogonal array $OA(vk,n,v,1)$ [22, Corollary 2.5]. A general orthogonal array $OA(N,n,v,t)$ of size $N$, degree $n$, order $v$, and strength $t$ is an $N$-set of $n$-tuples from $\{1,\ldots,v\}^{n}$ such that for any choice of $t$ coordinates, each $t$-tuple from $\{1,\ldots,v\}^{t}$ appears as a restriction of the $n$-tuples to the chosen coordinates exactly $\lambda$ times. This parameter is called the index of the orthogonal array and is given by $\lambda=N/v^{t}$. In our case the index is $k$, meaning that each element from $\{1,\ldots,v\}$ appears exactly $k$ times in every coordinate. This property is not sufficient; a characterization of $OA(vk,n,v,1)$’s representing $\mathcal{P}^{n}(v,k,\lambda)$-cubes is given in [22, Proposition 2.4].
@@
We performed this calculation for parameters $(7,3,1)$ and $(7,4,2)$, increasing the dimension until further extension is not possible. We relied on canonical labelings produced by nauty and Traces [26] to eliminate equivalent (partial) OA-representations of cubes. The results $\nu(7,3,1)=7$ and $\nu(7,4,2)=9$ are considerably smaller than the bound (2), which gives $28$ for $v=7$. Online versions of Table 1 and other tables in this paper are available on the web page
@@ 3. Autotopies
The claim follows from the observation that $(\pi_{x},\pi_{y})$ is an automorphism of the symmetric design $A=\Pi_{xy}(C)$. If $P_{x}$ and $P_{y}$ are the corresponding permutation matrices, then $A=P_{x}AP_{y}^{t}$ holds. By Lemma 2.1 we can write this as $P_{y}=A^{-1}P_{x}A$. Now $P_{x}$ and $P_{y}$ have the same trace, and this is the number of fixed points of $\pi_{x}$ and $\pi_{y}$. ∎
@@
On the other hand, autotopies of $\mathcal{C}$-cubes can have components with different numbers of fixed points. The $\mathcal{C}^{3}(7,3,1)$ cube of [21, Example 2.2] has an autotopy of order $7$ fixing $(0,0,7)$ points per component. In [21, Propositions 5.1 and 5.3] a total of $2396$ inequivalent $\mathcal{C}^{3}(16,6,2)$-cubes have been constructed. They have autotopies fixing $(0,0,f)$ points per component for $f=2,4,6,8,12,14,16$. This also provides counterexamples for bounds on the number of fixed points and other claims stated above.
@@
Table 3. $\mathcal{P}^{3}(16,6,2)$-cubes with an autotopy of order $8$ acting semiregularly.
Theorem 3.8.
In [21] and [22], constructions of $C^{3}(16,6,2)$ and $\mathcal{P}^{3}(16,6,2)$-cubes with prescribed autotopy groups $G$ were performed by a different approach. Essentially, all orthogonal arrays $OA(1536,3,16,2)$ and $OA(96,$ 3,16,1)3,16,1)3 , 16 , 1 ) invariant under $G$ were constructed and the ones not representing $\mathcal{C}$- and $\mathcal{P}$-cubes were discarded. Because of high proportions of “bad” OAs, rather large prescribed groups had to be used: $|G|\geq 512$ in [21] and $|G|\geq 18$ in [22]. We now have a more efficient approach for $\mathcal{P}$-cubes and can handle smaller groups. As an example we present the following result.
Proposition 3.9.
@@
A detailed statistics of the $1076$ cubes is given in Table 3. The cubes are divided according to the size of the full auto(para)topy group and the $2$-projections. The full autoparatopy group $\operatorname{Apar}(C)$ contains all combinations of isotopy and conjugation mapping $C$ onto itself. The group size is given as $(T,P)$, where $T=|\operatorname{Atop}(C)|$ and $P=|\operatorname{Apar}(C)|/T$. A question whether $\mathcal{P}^{3}(16,6,2)$-cubes with three non-isomorphic projections exist was raised in [22]. From the table we see that there are $152$ cubes with projections $(\mathcal{D}_{R},\mathcal{D}_{G},\mathcal{D}_{B})$ and the prescribed autotopy of order $8$. One example is shown in Figure 1, rendered in the ray tracing software POV-Ray [30]. Red, green, and blue light sources were placed so that the projections appear as shadows in the appropriate color.

### v2

@@ Abstract.
We study two kinds of generalizations of symmetric block designs to higher dimensions, the so-called $\mathcal{C}$-cubes and $\mathcal{P}$-cubes. For small parameters, all examples up to equivalence are determined by computer calculations. Known properties of automorphisms of symmetric designs are extended to autotopies of $\mathcal{P}$-cubes, while counterexamples are found for $\mathcal{C}$-cubes. An algorithm for the classification of $\mathcal{P}$-cubes with prescribed autotopy groups is developed and used to construct more examples. A bound on the dimension of difference sets for $\mathcal{P}$-cubes is proved and shown to be tight in elementary abelian groups. The construction is generalized to arbitrary groups by introducing regular sets of (anti)automorphisms.
@@ 1. Introduction
Here, $I$ is the identity matrix and $J$ is the all-ones matrix. An $n$-cube of $(v,k,\lambda)$ designs [21] is an $n$-dimensional $v\times\cdots\times v$ matrix such that its every $2$-section is an incidence matrix of a $(v,k,\lambda)$ design. Sections of dimension $2$ are submatrices obtained by fixing all but two coordinates. This generalization is a special case of W. de Launey’s proper $n$-dimensional transposable designs; see [5, Definitions 2.1, 2.6, 2.7, and Example 2.2]. The set of all $n$-cubes of $(v,k,\lambda)$ designs was denoted by $\mathcal{C}^{n}(v,k,\lambda)$ in [21]. We shall refer to them as $\mathcal{C}$-cubes.
The second generalization was introduced in [22] under the name $(v,k,\lambda)$ projection $n$-cubes. These are $n$-dimensional $v\times\cdots\times v$ matrices with $\{0,1\}$-entries such that every $2$-projection is an incidence matrix of a $(v,k,\lambda)$ design. If $C$ is an $n$-dimensional matrix and $1\leq x<y\leq n$, the projection $\Pi_{xy}(C)$ is the $2$-dimensional matrix with $(i_{x},i_{y})$-entry
@@
The sum is taken over all $n$-tuples $(i_{1},\ldots,i_{n})\in\{1,\ldots,v\}^{n}$ with fixed coordinates $i_{x}$ and $i_{y}$ in a field of characteristic 00. This definition was inspired by Room squares [10], which are generalized to $n$-dimensional Room cubes in an analogous way. In [22], the set of all $(v,k,\lambda)$ projection $n$-cubes was denoted by $\mathcal{P}^{n}(v,k,\lambda)$. We shall refer to them as $\mathcal{P}$-cubes. Two examples are shown in Figure 1, with three light-sources placed along the coordinate axes so that the projections appear as shadows. These and other pictures in the paper were rendered using the ray tracing software POV-Ray [30].
Figure 1. Cubes from $\mathcal{P}^{3}(7,4,2)$ and $\mathcal{P}^{3}(11,5,2)$.
The purpose of this paper is to study both types of cubes and to compare their properties. For dimension $n=2$, both are just incidence matrices of symmetric designs, but for $n\geq 3$ there are significant differences. For example, it was shown in [22, Theorem 2.9] that the dimension of $(v,k,\lambda)$ projection cubes with $k\geq 2$ is bounded by
$n\leq\frac{vk-1}{k-1}.$ (2)
@@
The organization of our paper is as follows. In Section 2, we recall the definitions of isotopy and equivalence of $n$-dimensional incidence cubes. We discuss the numbers of inequivalent cubes in $\mathcal{C}^{n}(v,k,\lambda)$ and $\mathcal{P}^{n}(v,k,\lambda)$ for $k=1$ and $k=2$. The main result in this section is Theorem 2.4, giving a complete enumeration up to equivalence of $\mathcal{P}$-cubes for parameters $(7,3,1)$, $(7,4,2)$, and all dimensions $n$. The proof is a computer calculation based on an algorithm that successively increases the dimension. The largest possible dimensions are $\nu(7,3,1)=7$ and $\nu(7,4,2)=9$.
@@
In Section 4, we study $\mathcal{P}^{n}(v,k,\lambda)$-cubes constructed from higher-dimensional difference sets. Theorem 4.1 gives a sharper bound than (2) on the dimension: $n\leq v$. Building on results from [22], we compute new values of $\mu_{G}(v,k,\lambda)$, the largest dimension of $(v,k,\lambda)$ difference sets in the group $G$. For elementary abelian groups $G$, Theorem 4.4 shows that the bound is tight, i.e. $\mu_{G}(v,k,\lambda)=v$ holds whenever difference sets exist. Theorem 4.7 generalizes the construction to arbitrary groups $G$ based on the notion of regular sets of (anti)automorphisms. A nice consequence is Corollary 4.8, showing that cyclic $(v,k,\lambda)$ difference sets extend at least to dimension $p$, where $p$ is the smallest prime divisor of $v$.
In the final Section 5, we put forward some observations based on data in Table 4. The table contains numbers of inequivalent $\mathcal{P}^{n}(v,k,\lambda)$-cubes constructed from difference sets. We have only partial explanations for apparent symmetries of the numbers and hope that the observations could lead to new theorems.
Some of our results have traditional formal proofs, while others are proved by computer calculations. There are numerous connections between the two types of results in both directions. In Section 3, formal results about autotopies enable computer classifications of $\mathcal{P}$-cubes with prescribed autotopy groups. In Section 4, computer classifications of higher-dimensional difference sets provide impetus for formal results, notably Theorems 4.4 and 4.7. Higher-dimensional incidence cubes are most easily explored by studying examples and performing experiments on a computer. Tools for examining $\mathcal{C}$- and $\mathcal{P}$-cubes of symmetric designs are available in the GAP [12] package Prescribed automorphism groups [20]. Plenty of examples are given in this paper, as well as in [21, 22].
@@ 2. Equivalence and classification
The classification problem is to determine the equivalence classes of $\mathcal{C}$- and $\mathcal{P}$-cubes for given parameters $(v,k,\lambda)$ and dimension $n$. We start with the degenerate case $k=1$. Incidence matrices of symmetric $(v,1,0)$ designs are permutation matrices of order $v$. They are all equivalent, since the rows and columns can be permuted to get the identity matrix $I$. This extends straightforwardly to $\mathcal{P}^{n}(v,1,0)$-cubes, which are all equivalent to
@@
On the other hand, the classification of $\mathcal{C}^{n}(v,1,0)$-cubes is a difficult problem. For $n=3$, they are in $1$-to-$1$ correspondence with latin squares $L=(\ell_{i_{1}i_{2}})$ of order $v$ by $C(i_{1},i_{2},i_{3})=[\ell_{i_{1}i_{2}}=i_{3}]$. The square bracket $[P]$ is the Iverson symbol, taking the value $1$ if $P$ is true, and 00 otherwise [19]. In Figure 2, representatives of the two main classes of latin squares of order $4$ and the corresponding $\mathcal{C}^{3}(4,1,0)$-cubes are shown. The number of main classes of latin squares has been determined by computer calculations up to $v=11$ [13, 25].
$\left(\begin{array}[]{llll}1&2&3&4\\ 2&3&4&1\\ 3&4&1&2\\ 4&1&2&3\end{array}\right)\kern 34.1433pt\left(\begin{array}[]{llll}1&2&3&4\\ 2&1&4&3\\ 3&4&1&2\\ 4&3&2&1\end{array}\right)$
Figure 2. Two $\mathcal{C}^{3}(4,1,0)$-cubes and the corresponding latin squares of order $4$.
Cubes in $\mathcal{C}^{n}(v,1,0)$ of arbitrary dimension $n$ are in $1$-to-$1$ correspondence with latin hypercubes of order $v$ and dimension $n-1$. There have been several definitions of latin hypercubes in the literature; the suitable definition for our purpose is used in [27]. A latin hypercube of order $v$ and dimension $n$ is a $v\times\cdots\times v$ matrix with entries from $\{1,\ldots,v\}$, such that every $1$-section contains every symbol exactly once. The set of all such objects is denoted by $\mathcal{H}^{n}_{v}$. In [27], classification is performed for $n=3$, $v\leq 6$ and $n=4,5$, $v\leq 5$.
@@
Cubes in $\mathcal{P}^{n}(3,2,1)$ have been classified directly in [22]. The result is reproduced in the first row of Table 1. We see that $\nu(3,2,1)=5$ and the bound (2) is tight in this case. The next feasible parameters are $(7,3,1)$. It is well known that the Fano plane is the unique $(7,3,1)$ design. In [22], two inequivalent $\mathcal{P}^{3}(7,3,1)$-cubes are presented as Examples 2.2 and 2.6. Here we present a full classification of $\mathcal{P}^{n}(7,3,1)$ and $\mathcal{P}^{n}(7,4,2)$-cubes.
@@
The bound (2) gives $\nu(7,3,1)\leq 10$ and is not tight for these parameters, but it is tight for $(7,4,2)$. The proof of Theorem 2.4 is a computer calculation that was carried out using the orthogonal array representation of $\mathcal{P}$-cubes. An orthogonal array $OA(N,n,v,t)$ of size $N$, degree $n$, order $v$, and strength $t$ is an $N$-set of $n$-tuples from $\{1,\ldots,v\}^{n}$ such that for any choice of $t$ coordinates, each $t$-tuple from $\{1,\ldots,v\}^{t}$ appears as a restriction of the $n$-tuples to the chosen coordinates exactly $\lambda$ times. This parameter is called the index of the orthogonal array and is given by $\lambda=N/v^{t}$.
A function $C:\{1,\ldots,v\}^{n}\to\{0,1\}$ is the characteristic function of a set of $n$-tuples
@@
If $C\in\mathcal{P}^{n}(v,k,\lambda)$, then $\overline{C}$ is an orthogonal array $OA(vk,n,v,1)$ of index $k$ [22, Corollary 2.5]. This means that each element from $\{1,\ldots,v\}$ appears exactly $k$ times in every coordinate. This property is not sufficient; a characterization of $OA(vk,n,v,1)$’s representing $\mathcal{P}^{n}(v,k,\lambda)$-cubes is given in [22, Proposition 2.4]. In the proof of [22, Theorem 2.9] it was also shown that $\overline{C}$ is a distance-invariant non-linear $v$-ary code of length $n$, in which only the distances $\{n-1,n\}$ occur.
@@
We performed this calculation for parameters $(7,3,1)$ and $(7,4,2)$, increasing the dimension until further extension is not possible. We relied on canonical labelings produced by nauty and Traces [26] to eliminate equivalent (partial) OA-representations of cubes. Online versions of Table 1 and other tables in this paper are available on the web page
@@ 3. Autotopies
The claim follows from the observation that $(\pi_{x},\pi_{y})$ is an automorphism of the symmetric design $A=\Pi_{xy}(C)$, for all $1\leq x<y\leq n$. If $P_{x}$ and $P_{y}$ are the corresponding permutation matrices, then $A=P_{x}AP_{y}^{t}$ holds. By Lemma 2.1 we can write this as $P_{y}=A^{-1}P_{x}A$. Now $P_{x}$ and $P_{y}$ have the same trace, and this is the number of fixed points of $\pi_{x}$ and $\pi_{y}$. ∎
@@
On the other hand, autotopies of $\mathcal{C}$-cubes can have components with different numbers of fixed points. The $\mathcal{C}^{3}(7,3,1)$ cube of [21, Example 2.2] has an autotopy of order $7$ with two components fixing no points, and the third component fixing $7$ points. In [21, Propositions 5.1 and 5.3] a total of $2396$ inequivalent $\mathcal{C}^{3}(16,6,2)$-cubes have been constructed. They have autotopies with two components fixing no points, and the third component fixing $f=2$, $4$, $6$, $8$, $12$, $14$, or $16$ points. This also provides counterexamples for bounds on the number of fixed points and other claims stated above.
@@
Theorem 3.8.
In [21] and [22], constructions of $C^{3}(16,6,2)$ and $\mathcal{P}^{3}(16,6,2)$-cubes with prescribed autotopy groups $G$ were performed by a different approach. Essentially, all orthogonal arrays $OA(1536,3,16,2)$ and $OA(96,$ 3,16,1)3,16,1)3 , 16 , 1 ) invariant under $G$ were constructed and the ones not representing $\mathcal{C}$- and $\mathcal{P}$-cubes were discarded. Because of high proportions of “bad” OAs, rather large prescribed groups had to be used: $|G|\geq 512$ in [21] and $|G|\geq 18$ in [22]. We now have a more efficient approach for $\mathcal{P}$-cubes and can handle smaller groups. As an example, we present the following result.
Proposition 3.9.
@@
Table 3. $\mathcal{P}^{3}(16,6,2)$-cubes with an autotopy of order $8$ acting semiregularly. Table of the statistics of the $1076$ cubes in Proposition 3.9.
@@
A detailed statistics of the $1076$ cubes is given in Table 3. The cubes are divided according to the size of the full auto(para)topy group and the $2$-projections. The full autoparatopy group $\operatorname{Apar}(C)$ contains all combinations of isotopy and conjugation mapping $C$ onto itself. The group size is given as $(T,P)$, where $T=|\operatorname{Atop}(C)|$ and $P=|\operatorname{Apar}(C)|/T$. A question whether $\mathcal{P}^{3}(16,6,2)$-cubes with three non-isomorphic projections exist was raised in [22]. From the table, we see that there are $152$ cubes with projections $(\mathcal{D}_{R},\mathcal{D}_{G},\mathcal{D}_{B})$ and the prescribed autotopy of order $8$. One example is shown in Figure 3, with red, green, and blue light sources so that the projections are in the appropriate color.